OpenStudy (please.help.me):

√2/2 csc x − 1 = 0 Solve the equation. (Enter your answers as a comma-separated list. Use n as an integer constant. Enter your response in radians.)

OpenStudy (please.help.me):

\[\frac{ \sqrt{2} }{ 2 } \csc x-1=0\]

jhonyy9 (jhonyy9):

hope you know that csc x = 1/sin x yes ? so rewrite this equation and will get sqrt2 1 ----- * ----- -1 = 0 add to both sides 1 and will get 2 sin x sqrt2 1 ---- * ----- = 1 multiply both sides by sin x and will get 2 sin x sqrt2 ----- = sin x 2 but we know that in case of sin x = sqrt2 /2 => x = arcsin(sqrt2 /2) what mean that x = 45 degree hope helped